OpenStudy (anonymous):

Someone please be ever so kind with this simplifying quotients problem(its just one fraction)

OpenStudy (anonymous):

\[\frac{ 2\sqrt{3} }{ \sqrt{6} }\]

OpenStudy (anonymous):

\[\sqrt{6}=\sqrt{3*2}\]

OpenStudy (anonymous):

That's all?

OpenStudy (anonymous):

\[\sqrt{3*2}=\sqrt{3}*\sqrt{2}\]

OpenStudy (anonymous):

the final ans is \[\frac{ 2 }{ \sqrt{2} }\]